OpenStudy (anonymous):

Hasan buys two kinds of cloth materials for school uniform shirt materials that costs him Rs.50 per metre and trouser material that costs him Rs.90 per metre.for every 3 metre of shirt material he buys 2 metre of the trouser material.He sells the material of 12% and 10% profit respectively.His total suits is Rs.36660.How mu trouser material he buy?

OpenStudy (anonymous):

i go the ans as

OpenStudy (anonymous):

shirt material = Rs 50 per metre profit=12% sp =(50*12/100) +50=56 trouser material = Rs 90 per metre. profit=10% sp =(90*10/100) +90=99 For every 2 meters of the trouser material he buys 3 metres of the shirt material. = (56*3)+(99*20) =366 let the number of sets he sold be y.

OpenStudy (anonymous):

is my ans is correct

OpenStudy (anonymous):

I believe so that is the same answer i got. because when you let the number of sets be y: total sell = 366*y=36600 so,y=100 every set contains 2 meters of trouser material. so Hasan sold 2*100=200 meters.

OpenStudy (anonymous):

@maariee Thank You:))